(no subject)

John Shields (jshields@ksera)
Thu, 24 Jun 1999 07:07:12 -0500

Shift Summary for Owl shift, 06/24/99

On Shift: J. Shields, J. Ladue

Hours of beam delivered at all intensities: 6.8
Hours of beam delivered at nominal intensity: 6.8
Hours of 99 data (special runs not included): 0.0
Hours of special runs (laser scans, muon runs...): 0.0 hrs

Accounting of lost time:

briefly lost beam, due to booster supply problem

========================================================================
Runs ended this shift:

Tapes brought to FCC this shift: 0

======================================================================
Accesses during shift: none

=====================================================================
Overview of Shift:

quiet, boring, and pretty much thoroughly uneventful. On the bright side,
we got a fair amount of beam and labeled lots of tapes.

=====================================================================

Description of Shift:

run spill time comment
------------------------------------------------------------------------
13706 00:46 disk usage on ktev4c exceeds 94%. We compressed
data files: sw_sdaq.dbs_Jun14 and 15,21,22.
Usage --> 84%. e-mailed Ashkan and Theo.

231 01:28 ktevana error: Missing FFFF word, stream 5 RWP 9
missed spill 232 (only) while running kfriend.
When we looked in the log for this error, the word
order appeared to be screwed up.

249 02:03 ktevana error: Missing FFFF word, etc (same as before)
ran kfriend and restarted triggers.

303 03:01 Scalar:Live Time reading out zero error shows up.
stopped/started triggers and kfriended.
Lost spill 303 and 304 as a result.

05:30 labeled 20 QKC tapes

05:45 Flushed with their success at labelling the QKC tapes,
the intrepid shift crew dares to tempt fate and
courageously decides to *also* label 20 QKK tapes.

05:50 Csi 2589 has comparater problems. The frequency
is now at 13, whereas it was only at 6 at 03:20.

440 06:22 Beam goes away. No explanation yet.

06:24 Message from beams division: Booster supply problems
estimate is 10 minutes.

460 06:30 Miraculously, beam is restored-- *ahead* of the
estimate! Go figure.

07:00 just as we pass the torch to the day shift, beam
is lost and won't be restored until 12 noon.

========================================================================
Follow up Notes on Previous Shift

========================================================================
Items to Follow up on Next Shift